# Sharing is CAIRing: Characterizing Principles and Assessing Properties of Universal Privacy Evaluation for Synthetic Tabular Data

### v1

@@ 3 Method
Table 1: Rubric consisting of four categories (Comparability, Applicability, Interpretability, Representativeness), each with four dimensions. Scoring is performed on a scale of 1-4 with 0.5 decimals allowed when a metric falls between two scores. The final result is the unweighted mean of the 16 scores.
@@
Scale
- Completely invariant to the unit scale of attributes.
@@
- The metric has a clearly defined upper and lower bound.
- The metric has an upper and lower bound.
@@
- The metric has either an upper or lower bound but not both.
- The metric has no upper or lower bounds.
@@
- Mixed data types have little to no influence on the analysis and do not significantly affect results or conclusions.
- Mixed data types introduce mild inaccuracies in the output.
@@
Cross-domain
relevance
- The metric shows comprehensive relevance across diverse domains.
@@
- The metric shows moderate cross-domain relevance, but may not cover all domains equally or comprehensively.
- The metric has limited relevance across different domains.
@@
Heterogeneity
- Applicable to most tabular data types.
@@
Diverse generation methods
- Applicable to all synthetic data sets regardless of generation method.
@@
- Is only defined for synthetic data generated using a single family of generative methods, e.g., GANs or CARTs.
- Can only be applied to synthetic data generated with one specific generative method.
Practicality
- The required runtime is reasonable. $\mathcal{O}(n^{2})$.
- The metric balances accuracy and runtime, but further improvements to efficiency can be accomplished. $\mathcal{O}(mn^{2})$ where $1<m\ll n$.
- The metric is computationally expensive. $\mathcal{O}(n^{3})$.
- Requires substantial computational resources. $\Omega(n^{3}).$
Implementation
- Easy to implement and/or an implementation is supplied.
@@
Explainability
- It is easy to communicate what the metric measures.
@@
Understandability
- The output of the metric is easy to understand for a layperson without the need for explanations.
- The output of the metric can be understood by a layperson, but might require some explanation.
- Understanding the output of the metric requires some background knowledge and is somewhat difficult to communicate to a layperson.
@@
Visualization
- The metric can easily be visualized both in isolation and compared to the results of other datasets.
@@
Granularity
- The metric offers fine-grained granularity that allows for unambiguous differentiation between privacy levels.
@@
Anomalies
- The sensitivity of all rare observations are taken into account.
@@
Coverage
- The privacy result is equally accurate for all records.
- The privacy result is approximately equally accurate for all records without any systematic bias.
- The privacy result is representative for some proper subsets of the data with a systematic bias.
- The privacy result only represents a small proper subset of the data.
Reproducibility
- The metric always returns the same value for the same input data.
- The metric returns the same value for the same input data with a probability close to 1 when applied by different people.
- The metric returns almost the same value for the same input data with some probability, but the variance is somewhat consistent when applied by different people.
@@
Precision
- The metric does not have systematic over- or underestimation of the privacy level.
- The metric slightly underestimates the privacy level but never overestimates it.
- The metric either over- or underestimates the privacy level, but not systematically.
- The metric significantly over- or underestimates the privacy level, systematically.
4 The CAIR Principles
The CAIR principles for good privacy evaluation for synthetic tabular data consist of a range of criteria that all aim to describe the ideal privacy measure. The rubric in Table 1 is limited in description depth, but should be sufficient to score any privacy metric without additional information. However, more detailed information on the dimensions is necessary to adequately argue for their inclusion.
4.1 Comparability
One of the major arguments for having standardized privacy evaluation metrics is the ability to compare privacy results between models, datasets, and projects. Comparisons enable users to select the best-performing models and measure them against the results of others, thus driving the synthetic data generation research area in a positive direction.
C1 Scale
Comparisons of privacy levels between generation processes rely on the privacy measure being on the same scale. If the scale is not identical, the resulting evaluation values do not correspond to each other, making any attempt at comparing the results futile. An argument can be made that users should normalize the data before computing the evaluation results. However, considering that synthetically generated data are primarily useful on the same scale as the true data, it is unlikely that all users will report the results on normalized data. In addition, the scale of the data can skew the results of a measure even if the metric has well-defined bounds. Consider the function $f(\mathcal{T},\mathcal{S})=\sigma(\lVert\mathcal{T}_{i}-\mathcal{S}_{j}\rVert_{2})$, where $\mathcal{T}$ and $\mathcal{S}$ are true and synthetic data respectively and $\sigma$ is the sigmoid function; then if an attribute, $a_{0}\in(0,10^{6})$ and another attribute $a_{1}\in(0,1)$, the influence $a_{0}$ has on the function will be substantially greater than that of $a_{1}$ even though the final result is in $(0,1)$.
C2 Metric Bounds
C3 Data Type Agnostic
Heterogeneous data types are prevalent in tabular data [14]. As such, it is crucial that the heterogeneity of the data has a limited effect on the output of the metric. Ideally, the metric should be invariant to any tabular data type, but a remediation can be to perform appropriate pre-processing. If pre-processing cannot mitigate adverse effects of the mixed data types, further analysis should be conducted to identify whether a systematic bias exists towards any specific subsets of data types.
C4 Cross-Domain Relevance
@@ 4.2 Applicability
The success of a universal privacy evaluation metric is highly dependent on the applicability of the metric. A universal metric should be applicable to any synthetically generated tabular data, regardless of how it was generated and regardless of the data types. Furthermore, it should be realistic to compute the metric with available information and computational resources.
@@
The heterogeneous nature of tabular data can pose a challenge when evaluating privacy. Many evaluation methods rely on a measure of similarity between synthetic and true records. However, many popular similarity functions, such as Euclidean and Hamming distance, perform poorly with more than one specific data type. Commonly, authors declare that any similarity function can be applied and rely on the reader to implement the metric with necessary adjustments [20, 4]. Proposing metrics with vague definitions restricts the ability to compare results across various implementations accurately. Accordingly, a good privacy evaluation metric considers how data with heterogeneous data types should be handled.
@@
A3 Practicality
Computational efficiency is highly relevant when considering how realistic the method is to use. Considering that many evaluation methods are based on pairwise distances [11, 14], the runtime should ideally be no more than $\mathcal{O}(n^{2})$ where $\mathcal{T},\mathcal{S}$ are the true and syntehtic data, respectively, and $n=\max(|\mathcal{T}|,|\mathcal{S}|)$.
@@
A privacy evaluation method is only useful if it is realistic to implement and compute. Ideally, a well-documented implementation of the evaluation metric should be provided whenever a new method is proposed. At the very least, the method should be well documented to allow for easy implementation by the reader and to avoid any ambiguity. Consequently, a measure will gain a low implementation score if it contains many complex elements that require extensive domain knowledge to implement appropriately. Furthermore, the data required to compute the metric should be readily accessible, without external data being required. We distinguish between internal and external information: Internal information is contained in the datasets for which the metric is computed, and external information is gathered elsewhere. An example of this property is illustrated in Section 5.
@@ 4.3 Interpretability
An important reason for ensuring the privacy level in synthetic data is to enable data sharing of otherwise sensitive data. Regulatory agencies must recognize that adequate privacy levels are provided to accomplish that goal. Consequently, it is crucial that privacy levels can be accurately communicated to relevant stakeholders, be they data regulators or organizational managers. The Interpretability category aims to put emphasis on this aspect of creating universal privacy evaluation metrics.
@@
What is measured and how it evaluates privacy should be easy to explain to any stakeholder. Using an explainable metric increases the transparency and, thereby, improves the confidence in the metric. Additionally, explainability increases the probability that the privacy provided by synthetic data is recognized and accepted by any relevant stakeholder. Accordingly, a good privacy measure is made up of elements that are simple to relay to laypeople.
@@
Granularity refers to a user’s ability to unambiguously differentiate between privacy levels. The most basic form of fine-grained differentiation is to ensure that the privacy measure follows a monotonic function. Any non-monotonic function can confuse a user, especially laypeople. Furthermore, a linear function is preferred as the privacy levels change at a constant rate as a function of the underlying computation. Consequently, any privacy metric that forces well-defined bounds by transformation, for example, by using a sigmoid or hyperbolic tangent function, will be slightly penalized. In addition, the level of granularity decreases if the output of the metric consists of more than a single value. Keeping track of more than one value means that the user must understand and interpret the individual outputs and how they change in connection. Lastly, the privacy level should be understandable in isolation. The results of some measures may require comparisons with the results of other datasets or generation methods to accurately differentiate between privacy levels. Such metrics should be penalized for this need.
@@ 4.4 Representativeness
The last category scores the ability of the privacy metric to accurately measure privacy levels. If the metric cannot capture potential privacy violations, the metric is not usable as a privacy metric.
@@
Anomalous records typically have a higher risk of being subject to privacy violations [20]. Inliers typically have the benefit of “blending in” with other common observations, ensuring that even highly similar synthetic records provide sufficient privacy levels. Outliers do not enjoy the same security. Synthetic records in the immediate proximity of a true record are likely to be uniquely linked to each other, constituting a privacy violation. Accordingly, anomalous observations should be given appropriate considerations when evaluating the overall level of privacy. Considering that many methods use some form of similarity measure, methods can be constructed to include weighted versions of the similarity function with the weights representing the degree to which individual records are outliers. The outcome of this is that records close to an outlier are made to be more similar to the outlier than would be the case if an unweighted distance function were used.
@@ 5 Application of CAIR
Figure 2: Radar plots of CAIR scores for the selected privacy evaluation metrics. Dimensions are indexed according to the system in Section 4 and separated into their respective categories denoted by the legend.
A benefit of having a scoring system to assess evaluation metrics is the ability to score existing metrics and rank them accordingly. As an illustrative example, a selection of tabular evaluation metrics is selected and scored with the results presented in Figure 2 (details on the scoring in Appendix A).
@@
Distance to the closest record (DCR) is a simple metric where the output is typically the mean or median minimum distance from a synthetic record to a true one [9]. DCR can be used as a metric on its own or it can be extended to membership inference.
Threshold-based membership inference attacks (T. MIA) are attacks in which an adversary attempts to infer whether a known individual has been used to generate synthetic data [19]. Evaluation can be carried out in many ways, but a naïve approach is to use DCR for each record and to predict it as a member if $\text{DCR}\leq\tau$ where $\tau$ is a fixed threshold.
@@
Attribute disclosure attacks (Attr. disclosure) are similar to MIAs, but instead of inferring membership, the adversary attempts to predict unknown attribute values [13]. A $k$-nearest neighbor algorithm can be used, and the majority vote predicts the univariate values for $k>1$. The attack is typically performed on various levels of prior adversarial knowledge and the output is all of the results or aggregate statistics.
@@
$\epsilon$-identifiability ($\epsilon$-Ident.) takes a different approach and reports the proportion of synthetic records that are “too similar” to the true data points [20]. Let $d$ be the minimum weighted distance within the true data, let $\hat{d}$ be the weighted distance between the true and a specific synthetic record, and let $n$ be the number of synthetic records; then $\epsilon\text{-iden.}=\frac{1}{n}[\mathbb{I}(\hat{d}<d)]$, where $\mathbb{I}$ is the identity function.
@@
Identity disclosure risk (IDR) incorporates functions like $\epsilon$-identifiability, but extends them with additional information, some of which are not readily available [4]. The function takes the inverse size of the equivalence class and multiplies by: 1) an indicator variable, such as $\epsilon$-identifiability, 2) an error correction term from the literature, and 3) a binary variable indicating whether the adversary learns something new. The equivalence class from both the sample and the true population is required.

### v2

@@ 4.1 Comparability
Table 1: Rubric consisting of four categories (Comparability, Applicability, Interpretability, Representativeness), each with four dimensions. Scoring is performed on a scale of 1-4 with 0.5 decimals allowed when a metric falls between two scores.
@@
Scale - Completely invariant to the unit scale of attributes.
@@
- The metric has a clearly defined upper and lower bound. - The metric has an upper and lower bound.
@@
- The metric has either an upper or lower bound but not both. - The metric has no upper or lower bounds.
@@
- Mixed data types have little to no influence on the analysis and do not significantly affect results or conclusions. - Mixed data types introduce mild inaccuracies in the output.
@@
Cross-domain relevance - The metric shows comprehensive relevance across diverse domains.
@@
- The metric shows moderate cross-domain relevance, but may not cover all domains equally or comprehensively. - The metric has limited relevance across different domains.
@@
Heterogeneity - Applicable to most tabular data types.
@@
Diverse generation methods - Applicable to all synthetic data sets regardless of generation method.
@@
- Is only defined for synthetic data generated using a single family of generative methods, e.g., GANs or CARTs. - Can only be applied to synthetic data generated with one specific generative method.
Performance - The required runtime is reasonable. $\mathcal{O}(n^{2})$. - The metric balances accuracy and runtime, but further improvements to efficiency can be accomplished. $\mathcal{O}(mn^{2})$ where $1<m\ll n$. - The metric is computationally expensive. $\mathcal{O}(n^{3})$. - Requires substantial computational resources. $\Omega(n^{3}).$
Implementation - Easy to implement and/or an implementation is supplied.
@@
Explainability - It is easy to communicate what the metric measures.
@@
Understandability - The output of the metric is easy to understand for a layperson without the need for explanations. - The output of the metric can be understood by a layperson, but might require some explanation. - Understanding the output of the metric requires some background knowledge and is somewhat difficult to communicate to a layperson.
@@
Visualization - The metric can easily be visualized both in isolation and compared to the results of other datasets.
@@
Granularity - The metric offers fine-grained granularity that allows for unambiguous differentiation between privacy levels.
@@
Anomalies - The sensitivity of all rare observations are taken into account.
@@
Coverage - The privacy result is equally accurate for all records. - The privacy result is approximately equally accurate for all records without any systematic bias. - The privacy result is representative for some proper subsets of the data with a systematic bias. - The privacy result only represents a small proper subset of the data.
Reproducibility - The metric always returns the same value for the same input data. - The metric returns the same value for the same input data with a probability close to 1 when applied by different people. - The metric returns almost the same value for the same input data with some probability, but the variance is somewhat consistent when applied by different people.
@@
Precision - The metric does not have systematic over- or underestimation of the privacy level. - The metric slightly underestimates the privacy level but never overestimates it. - The metric either over- or underestimates the privacy level, but not systematically. - The metric significantly over- or underestimates the privacy level, systematically.
@@ 4.2 Applicability
The efficacy of a universal privacy evaluation metric is highly dependent on its applicability across diverse datasets and data generation methods. A universally applicable metric must have the ability to assess the privacy-preserving capabilities of synthetically generated tabular data comprehensively, irrespective of the generation techniques employed or the data characteristics. Furthermore, it should be feasible to compute the metric using readily available information and computational resources. This aspect of applicability ensures that the metric remains practical and accessible, allowing stakeholders to evaluate privacy effectively across a broad range of synthetic datasets without significant computational challenges or data-specific constraints.
@@
The heterogeneous nature of tabular data can pose a challenge when evaluating privacy. Many evaluation methods rely on a measure of similarity between synthetic and true records. However, many popular similarity functions, such as Euclidean and Hamming distance, perform poorly with more than one specific data type. Commonly, authors declare that any similarity function can be applied and rely on the reader to implement the metric with necessary adjustments (El Emam et al.,, 2020; Yoon et al.,, 2020). Proposing metrics with vague definitions restricts the ability to compare results across various implementations accurately. Accordingly, a good privacy evaluation metric considers how data with heterogeneous data types should be handled.
@@
A3 Performance
Computational efficiency is crucial when evaluating the practicality of a method, especially when considering the scalability to larger datasets. A method may provide accurate or granular results, but if it is computationally expensive, its applicability diminishes. Many evaluation methods for synthetic data are based on pairwise distances, which involves comparing every point in the true dataset $\mathcal{T}$ with every point in the synthetic dataset $\mathcal{S}$ (Hernandez et al.,, 2022; Murtaza et al.,, 2023). In this context, the performance of a metric should ideally not exceed $\mathcal{O}(n^{2})$ where $n=\max(|\mathcal{T}|,|\mathcal{S}|)$. While $\mathcal{O}(n^{2})$ is not the most efficient time complexity, it is the worst case for pairwise comparisons, thus defining a natural threshold for performance.
@@
A privacy evaluation method is only useful if it is realistic to implement and compute. Ideally, a well-documented implementation of the evaluation metric should be provided whenever a new method is proposed. At the very least, the method should be well-documented to allow for easy, consistent implementation by the reader and to avoid any ambiguity. Consequently, a measure will gain a low implementation score if it contains many complex elements that require extensive domain knowledge to implement appropriately. Furthermore, the data required to compute the metric should be readily accessible, without external data being required. We distinguish between internal and external information: Internal information is contained in the datasets for which the metric is computed, and external information is gathered elsewhere. An example of this property is illustrated in the application of CAIR later in the paper.
External information includes, but is not limited to, dynamic statistics, additional datasets for comparisons or similar, and evaluation results beyond the dataset that is currently being evaluated.
@@ 4.3 Interpretability
A fundamental objective of preserving privacy in synthetic data is to enable data sharing without compromising individual privacy rights. Ultimately, regulatory bodies define what constitutes adequate privacy levels, making them pivotal stakeholders to privacy preservation. The Interpretability principle of CAIR governs stakeholders’ ability to comprehend how privacy levels are measured and, ultimately, how data privacy is established through data synthesis. Therefore, it is crucial that the privacy levels determined by evaluation metrics can be effectively communicated and understood by the relevant stakeholders, including data regulators and organizational decision-makers.
@@
What is measured and how it evaluates privacy should be easy to explain to any stakeholder. Using an explainable metric increases the transparency and, thereby, improves the confidence in the metric. Explainability can be understood in multiple ways including using easy-to-understand concepts. Relying on complex formulas and black-box models can decrease the explainability to non-technical stakeholders, thus creating barriers for adopting the metric in regulatory decisions. To mitigate the perceived complexity of an evaluation metric, clear formulations and intuitive descriptions can be provided with non-technical stakeholders in mind. A metric that includes elements that can only be understood with technical knowledge inhibits comprehension. Therefore, explainability increases the probability that the privacy provided by synthetic data is recognized and accepted by any relevant stakeholder.
@@
Granularity refers to a user’s ability to unambiguously differentiate between privacy levels. The most basic form of fine-grained differentiation is to ensure that the privacy measure follows a monotonic function. Any non-monotonic function can confuse a user, especially laypeople. Furthermore, a linear function is preferred as the privacy levels change at a constant rate as a function of the underlying computation. Consequently, any privacy metric that forces well-defined bounds by transformation, for example, by using a sigmoid or hyperbolic tangent function, will be slightly penalized. In addition, the level of granularity decreases if the output of the metric consists of more than a single value such as precision and recall for an adversarial membership inference attack (Sun et al.,, 2023). Keeping track of more than one value means that the user must understand and interpret the individual outputs and how they change in connection. Lastly, the privacy level should be understandable in isolation. The results of some measures may require comparisons with the results of other datasets or generation methods to accurately differentiate between privacy levels. Such metrics should be penalized for this need.
@@ 4.4 Representativeness
The concept of Representativeness captures the ability of privacy metrics to accurately gauge privacy levels of synthetic datasets. An effective metric must possess the capacity to comprehensively capture privacy violations and quantify the severity of these violations. Essentially, it should reflect the complexity of real-world privacy concerns while equally representing the privacy levels of all individuals. Failure to sufficiently capture privacy violations undermines the validity of the metric. Thus, the Representativeness principle is an essential part of assessing the efficacy of privacy evaluation metrics for synthetic data.
@@
Anomalous records typically have a higher risk of being subject to privacy violations (Yoon et al.,, 2020). Inliers typically have the benefit of “blending in” with other common observations, ensuring that even highly similar synthetic records provide sufficient privacy levels. Outliers do not enjoy the same security. Synthetic records in the immediate proximity of a true record are likely to be uniquely linked to each other, constituting a privacy violation. Accordingly, anomalous observations should be given appropriate considerations when evaluating the overall level of privacy. Considering that many methods use some form of similarity measure, methods can be constructed to include weighted versions of the similarity function with the weights representing the degree to which individual records are outliers. Weighting of anomalous records results in records close to an outlier seeming more similar to the outlier than is the case when using an unweighted distance function.
@@ 5 Application of CAIR
Figure 4 illustrates the workflow of the demonstration in this section. First, a diverse selection of privacy evaluation metrics is identified. Two independent evaluators score the metrics on the 16 dimensions of CAIR. Aggregating their scores according to Figure 2 allows for a dimension-wise representation for a comprehensive analysis. Lastly, the assessments are interpreted in two ways: internally and externally. Internally, the strengths and weaknesses of each metric are identified, allowing high-fidelity insights that can lead to future improvements development strategies. External evaluation of the metrics involves comparing the results between all the metrics, both in terms of the individual dimensions and the overall CAIR score that allows for a ranking of the metrics based on their universality.
The results are depicted in Figure 5 with the overall score and error being computed using Equations 2 and 3, respectively. The scores provided in this section serve as an illustrative example rather than a definitive assessment. This section covers how CAIR can be used to identify areas of improvement for metrics and to gauge what metrics best conform to the concept of universality. Furthermore, the resulting ranking reflects the versatility of the privacy evaluation metrics, rather than solely their ability to precisely gauge the level of privacy. For a comprehensive analysis of each metric’s performance, the dimension-wise representation in Figure 3(c) is selected as the appropriate level of granularity. Detailed information on the scorings is provided in A.
Figure 4: Workflow for the application setup, demonstrating how CAIR is used and interpreted. A selection of metrics is assessed by two independent evaluators. Their scores are combined to dimension-wise means. Interpreting the results are conducted in two parts: 1) internally where the dimension-wise scores are assessed and 2) externally, where the metric assessments are compared to each other. Metric acronyms: DCR (Distance to Closest Record); Attr. Dis. (Attribute Disclosure); T. MIA (Threshold-based Membership Inference Attack); $\epsilon$-ID ($\epsilon$-Identifiability); IDR (Identity Disclosure Risk).
5.1 Metrics for Evaluation
To ensure a comprehensive demonstration of the capabilities of CAIR, a diverse selection of metrics is required. Privacy metrics for synthetic tabular data can be divided into three groups: Membership inference attacks (MIA), Attribute disclosure Risk, and similarity-based measures (Murtaza et al.,, 2023; Ganev and Cristofaro,, 2023). For each of these categories, a wide variety of metrics have been proposed. For this demonstration, we select metrics from all three categories, ranging from simple naïve to more complex measures.
@@ 5.1 Metrics for Evaluation
Distance to the closest record (DCR) is a simple metric where the output is typically the mean or median minimum distance from a synthetic record to a true one (Guillaudeux et al.,, 2023). In the most basic scenario, the DCR is computed using Euclidean distance. DCR can be used as a metric on its own, or it can be extended to membership inference.
@@
Attribute disclosure attacks (Attr. disclosure) are similar to MIAs, but instead of inferring membership, the adversary attempts to predict unknown attribute values (Kaur et al.,, 2021). An adversary is assumed to have access to a dataset similar to the synthetic data but with at least one attribute missing. The adversary can use a $k$-nearest neighbor algorithm, and the majority vote predicts the univariate values for $k>1$. The attack is typically performed on various levels of prior adversarial knowledge and the output is all of the results or aggregate statistics. Similar to threshold-based membership inference, the reported adversarial performance is the mean precision and recall over all levels of prior knowledge.
Threshold-based membership inference attacks (T. MIA) are attacks in which an adversary attempts to infer whether a known individual has been part of the training data for the generative model (Yan et al.,, 2020). Evaluation can be carried out in many ways, but a naïve approach is to use DCR for each record and to predict it as a member if $\text{DCR}\leq\tau$ where $\tau$ is a fixed threshold. The reported metric can be any that evaluates the performance of the adversarial attack, but for this evaluation, the reported metrics are assumed to be precision and recall.
DOMIAS takes a different approach to MIAs by investigating differences in densities for the synthetic data and an adversarial dataset (van Breugel et al.,, 2023). The method is a black-box setting assuming that an adversary has access to synthetic data, $\mathcal{S}$, and a dataset, $\mathcal{A}$, derived from the same population as the original training data, thus potentially containing instances used to train the synthetic data generation model. By producing density estimates for both $\mathcal{S}$ and $\mathcal{A}$, the method assumes overfitting in areas with higher density in $\mathcal{S}$ compared to $\mathcal{A}$. If a record in $\mathcal{A}$ is located within an overfitted area, membership is inferred. The performance is typically reported as either AUROC or accuracy.
@@
$\epsilon$-identifiability ($\epsilon$-id) takes a different approach and reports the proportion of synthetic records that are “too similar” to the true data points (Yoon et al.,, 2020). Let $d$ be the minimum weighted distance within the true data, let $\hat{d}$ be the weighted distance between the true and a specific synthetic record, and let $n$ be the number of synthetic records; then $\epsilon\text{-id}=\frac{1}{n}[\mathbb{I}(\hat{d}<d)]$, where $\mathbb{I}$ is the identity function.
@@
Identity disclosure risk (IDR) incorporates functions like $\epsilon$-identifiability, but extends them with additional information, some of which are not readily available (El Emam et al.,, 2020). The function takes the inverse size of the equivalence class and multiplies by: 1) an indicator variable, such as $\epsilon$-identifiability, 2) an error correction term from the literature, and 3) a binary variable indicating whether the adversary learns something new. The equivalence class is required from both the sample and the true population.
This selection of metrics is naturally non-exhaustive, but the diversity allows for a comprehensive analysis of the capabilities of CAIR and its ability to cover a wide range of privacy metrics and notions. Numerous other metrics exist in the literature, and their omission here is solely for brevity.
5.2 Findings
Figure 5: Radar plots of CAIR scores for the selected privacy evaluation metrics. The dimensions are indexed according to the general category and respective number and further separated into the corresponding principle denoted by the legend.